OpenStudy (anonymous):

Two coils having equal resistance but different inductances are connected in series, then the time constant of series combination will ______ a. sum of time constants of individual coils b. average of time constant of individual coils c. geometric mean of the time constants of individual coils d. product of the time constants of individual coils... please help me with the answer and tell me how...

OpenStudy (kenljw):

The time constant of an Inductor is L/R, series resistors and inductors add. Therefore the new time constant would be (L1 + L2)/(R! + R2) = (L1 + L2)/(2R) = (L1/R +L2/R)/2 therefore b The above assumes no mutual inductance due to closeness.